### v1

# Grasp2Grasp: Vision-Based Dexterous Grasp Translation via Schrödinger Bridges

Abstract
We propose a new approach to vision-based dexterous grasp translation, which aims to transfer grasp intent across robotic hands with differing morphologies. Given a visual observation of a source hand grasping an object, our goal is to synthesize a functionally equivalent grasp for a target hand without requiring paired demonstrations or hand-specific simulations. We frame this problem as a stochastic transport between grasp distributions using the Schrödinger Bridge formalism. Our method learns to map between source and target latent grasp spaces via score and flow matching, conditioned on visual observations. To guide this translation, we introduce physics-informed cost functions that encode alignment in base pose, contact maps, wrench space, and manipulability. Experiments across diverse hand-object pairs demonstrate our approach generates stable, physically grounded grasps with strong generalization. This work enables semantic grasp transfer for heterogeneous manipulators and bridges vision-based grasping with probabilistic generative modeling.
00footnotetext: Buchanan contributed to this work in the capacity as an independent researcher. The views expressed (or the conclusions reached) are their own and do not necessarily represent the views of Lockheed Martin Corporation.
1 Introduction
Robotic grasping remains a central challenge in autonomous manipulation, especially in the context of dexterous, multi-fingered hands. Unlike parallel-jaw grippers or simple suction devices, dexterous hands offer significantly more control and flexibility but also bring high-dimensional configuration spaces [2, 15] and non-trivial contact dynamics [61]. These properties pose difficulties [49] in data collection, modeling, and learning robust grasp strategies. While recent works [42, 48, 85, 8, 93] in data-driven grasp generation have shown success in generating feasible grasps for specific hand-object pairs, they often rely on large-scale datasets [42, 81, 88, 29] curated through computationally expensive simulation or real-world trials. Moreover, such models typically operate in a hand-specific setting, which requires retraining or significant fine-tuning to adapt to new robotic hands [37] and poses a practical bottleneck in the rapidly evolving landscape of robotic hardware.
This lack of generalization across different robotic hands severely limits the applicability of learned manipulation policies in real-world systems [91]. In particular, with the emergence of increasingly diverse robotic platforms, it becomes infeasible to collect training data or train models from scratch for every possible hand. Instead, it is desirable to reuse grasp knowledge acquired from one hand to infer meaningful grasp strategies for another. However, the morphological differences between hands mean that naive transfer, such as copying joint angles or end-effector poses, might lead to physically invalid or unstable grasps [64, 54, 1]. This motivates the need for a framework that can translate the intent behind a demonstrated grasp from a source hand to a new target hand, taking into account both the underlying object geometry and the structural differences in hand kinematics.
This paper aims to address the problem of vision-based dexterous grasp translation. As shown in Fig. 1, given a visual observation of a source robotic hand grasping an object, we aim to generate a plausible grasp for a target hand with a different morphology. This setting arises naturally in scenarios such as human-to-robot imitation [67, 90], hardware substitution in manipulation pipelines [27], and learning from heterogeneous demonstrations [34]. Although reminiscent of teleoperation [33, 21] or behavior cloning [14, 91], our formulation diverges in a critical way: the goal is not to replicate the precise joint configuration or trajectory of the source hand.
Figure 1: (top) Comparison of morphology and scale across hands. (bottom) Object-agnostic methods often miss fine-grained contacts, leading to invalid grasps. Our method produces contact-consistent, stable grasps across diverse hand morphologies.
Instead, we seek to preserve the functional intent of the grasp, expressed through physical quantities such as contact locations, grasp wrenches, or stability measures, while accounting for the morphological differences between the source and target manipulators.
To this end, we cast the grasp translation task as a probabilistic transport problem between the distribution of grasps executed by the source hand and those executable by the target hand. We leverage the formalism of the Schrödinger Bridge (SB) [71, 41], which models the most likely stochastic process interpolating between two distributions under a reference dynamics. This perspective allows us to frame grasp translation as learning a distributional flow from observed source grasps to compatible target grasps, guided by a meaningful notion of grasp similarity and grounded in physics.
Our approach builds on a growing body of work [18, 84, 13, 87, 73, 43, 78] that frames generative modeling as a Schrödinger Bridge problem, which enables probabilistic transport between arbitrary data distributions under stochastic dynamics. In particular, recent developments have demonstrated how score-based [18, 87] and flow-based [78, 19, 79] objectives can be leveraged to learn such bridges without the need for simulating full stochastic processes, offering scalable and flexible tools for distribution alignment. We adapt and extend these ideas to the domain of dexterous grasp translation, where the source and target distributions represent physically meaningful grasps from different robotic hands. To this end, we develop a latent-space learning pipeline for distribution-level grasp translation conditioned on vision-based observations of the source hand. Within this framework, we design custom ground cost functions that capture task-relevant physical properties, such as contact patterns and grasp stability, rather than relying on generic Euclidean or geodesic distances. This allows us to generate grasps that are not only kinematically feasible for the target hand but also functionally aligned with the intent of the source grasp.
We summarize our contributions as follows:
  • We formulate vision-based dexterous grasp translation as a Schrödinger Bridge problem between source and target grasp distributions, capturing the stochastic nature of plausible grasps.
  • We introduce novel ground cost functions tailored to the grasping domain, enabling semantic and physically-informed translation across different hand morphologies.
  • We develop a latent-space learning pipeline based on score and flow matching, enabling simulation-free, distribution-level training on unpaired visual grasp observations.
  • We evaluate our method across diverse hand-object combinations, demonstrating that our approach generates functionally equivalent and physically robust grasps with strong generalization capabilities.
2 Related Work
Dexterous Grasp Generation has been a long-standing topic in robotics, aimed at enabling robotic hands with multiple degrees of freedom (DoFs) to securely and effectively manipulate diverse objects. Early approaches [58, 20, 26, 7, 63, 5] relied on analytical models of hand kinematics and contact mechanics [56, 6] to compute force-closure [63, 60, 58, 45, 81] or form-closure [80, 55] grasps, which are often constrained by assumptions of perfect information and simple object geometries [56, 6]. The advent of vision-based data-driven techniques has led to significant advances [62, 92], particularly through direct regression [70, 44] or generative modeling techniques [31, 38, 30] such as VAEs [42, 57, 37], GANs [51, 57, 50, 16], and diffusion models [89, 32, 82, 48]. These generative approaches enable the synthesis of diverse grasp configurations conditioned on visual inputs such as RGB-D images [35, 36] or segmented point clouds [89, 32, 48], allowing for more flexible and scalable grasp generation. For example, GenDexGrasp [42] introduces a hand-agnostic grasping algorithm trained on the MultiDex dataset [42], which leverages contact maps as intermediate representations to efficiently generate diverse and plausible grasping poses transferable among various multi-fingered robotic hands. Similarly, UGG [48] presents a unified diffusion-based dexterous grasp generation model that operates within object point cloud and hand parameter spaces. While these generative approaches have advanced the field, they often remain tailored to specific robotic hands and lack mechanisms for transferring grasp knowledge across different hand morphologies. Our work addresses these limitations by developing a grasp translation framework that operates at the distributional level, enabling the reuse of grasp knowledge across different robotic platforms through a physically meaningful transport mechanism.
Grasp Transfer is closely related to teleoperation and imitation learning [3], where the objective is to map motions or actions from a source agent (e.g., a human or a robot) to a target robot. Teleoperation methods [72, 74, 95, 33, 21, 66] often aim for one-to-one replication of joint trajectories or end-effector poses, which is generally effective when the source and target share similar kinematics. However, such approaches struggle when transferring grasps between manipulators with significantly different morphologies, as direct replication can lead to infeasible or suboptimal grasps [64, 54, 1]. More recent work, such as RobotFingerPrint [37], addresses this challenge by mapping grippers into a shared Unified Gripper Coordinate Space (UGCS) based on spherical coordinates to facilitate grasp transfer. While promising, this approach requires simulation-based preprocessing and detailed gripper models, which impose a strong prior that limits flexibility. Moreover, the UGCS approach focuses on point correspondences instead of explicitly preserving functional grasp intent. In contrast, our formulation aims to translate grasps in a distributional sense grounded in physical properties. It requires no hand-specific simulation preprocessing, which enables broader applicability and more semantically meaningful grasp transfer.
Optimal Transport (OT) and Schrödinger Bridges (SB) have recently emerged as powerful frameworks in the machine learning community for modeling distributional transformations. OT [69, 86, 65, 75] provides a principled way to compute mappings between probability distributions that minimize a cost function, typically grounded in geometric distances. While classical OT has found applications in tasks like domain adaptation [17, 24] and shape matching [22, 40, 77], it often requires solving computationally intensive optimization problems [75, 59]. Schrödinger Bridges [71, 41] extend OT by introducing a stochastic reference process, which enables more flexible and entropically regularized transport. This has proven particularly useful in generative modeling [18, 84, 13, 87, 73, 43, 78], where SBs offer a probabilistic interpolation between data distributions. Recent advances in simulation-free training methods [46, 78, 79] have made these techniques more scalable by bypassing the need to simulate stochastic processes during training. In robotics, these frameworks have been explored for motion planning [39], distribution alignment [76], and dynamical system modeling [9]. However, their application to grasp synthesis and translation remains underexplored. Our work leverages SBs not just for smooth transport between distributions, but for preserving task-specific physical properties of grasps across different hand morphologies, pushing the boundaries of OT/SB utility in robotic manipulation.
3 Preliminaries
3.1 Schrödinger Bridge Formulation
The Schrödinger Bridge (SB) problem [71, 41] provides a principled framework for defining the most likely stochastic evolution between two probability distributions under a reference process. Given source and target distributions $q_{0}(x)$ and $q_{1}(x)$, and a reference stochastic process $\mathbb{Q}$, the SB seeks a process $\mathbb{P}$ that minimizes the relative entropy (Kullback-Leibler divergence) with respect to $\mathbb{Q}$, while ensuring that $\mathbb{P}$ matches the marginals $q_{0}$ and $q_{1}$ at times 0 and 1, respectively:
$\mathbb{P}^{*}=\operatorname*{arg\,min}_{\mathbb{P}:p_{0}=q_{0},p_{1}=q_{1}}\mathrm{KL}(\mathbb{P}||\mathbb{Q}).$ (1)
Typically, $\mathbb{Q}$ is chosen as a Brownian motion with a diffusion rate $\sigma$, leading to the so-called diffusion Schrödinger Bridge [18, 10]. The optimal process $\mathbb{P}^{*}$ can be characterized as a mixture of Brownian bridges weighted by an entropically regularized optimal transport (OT) plan. Specifically, the entropic OT plan $\pi_{\varepsilon}^{*}(q_{0},q_{1})$ minimizes a regularized cost:
$\pi_{\varepsilon}^{*}=\operatorname*{arg\,min}_{\pi\in U(q_{0},q_{1})}\int d(x_{0},x_{1})^{2}d\pi(x_{0},x_{1})+\varepsilon\mathrm{KL}(\pi||q_{0}\otimes q_{1}),$ (2)
where $U(q_{0},q_{1})$ denotes the set of couplings between $q_{0}$ and $q_{1}$, and $d(\cdot,\cdot)$ is a ground cost. Föllmer [28] connects this formulation to the SB problem: if the reference process $\mathbb{Q}$ is Brownian motion, the SB is uniquely characterized by a mixture of Brownian bridges drawn according to the entropic OT plan. This insight allows the complex dynamic optimization of SB to be reduced to computations over static distributions, making it more tractable in high-dimensional settings.
3.2 Score-based and Flow-Matching Methods
To model the stochastic dynamics of the SB, we consider Itô SDEs of the form:
$dx_{t}=u_{t}(x_{t})dt+g(t)dw_{t},$ (3)
where $u_{t}(x)$ is a drift vector field, $g(t)$ is a diffusion schedule, and $w_{t}$ is standard Brownian motion. The marginal densities $p_{t}(x)$ induced by this SDE evolve according to the Fokker-Planck equation [68]:
$\frac{\partial p_{t}}{\partial t}=-\nabla\cdot(p_{t}u_{t})+\frac{g(t)^{2}}{2}\Delta p_{t},$ (4)
where $\Delta p_{t}=\nabla\cdot(\nabla p_{t})$ is the Laplacian. An equivalent deterministic process can be described via the probability flow ODE:
$dx_{t}=\underbrace{\left(u_{t}(x_{t})-\frac{g(t)^{2}}{2}\nabla\log p_{t}(x_{t})\right)}_{u_{t}^{\circ}(x)}dt,$ (5)
which shares the same time marginals $p_{t}$ as the original SDE. The drift term $u_{t}(x)$ can be recovered from the flow drift $u_{t}^{\circ}(x)$ and score $\nabla\log p_{t}(x)$ as $u_{t}(x)=u_{t}^{\circ}(x)+\frac{g(t)^{2}}{2}\nabla\log p_{t}(x)$.
Learning such stochastic dynamics from data typically involves estimating both $u_{t}^{\circ}(x)$ and the score $\nabla\log p_{t}(x)$. In our work, we follow the [SF]2[78] formulation, which proposes to learn these quantities directly by regressing neural networks $v_{\theta}(t,x)$ and $s_{\theta}(t,x)$ to approximate them. In practice, the marginals $p_{t}$ are often not directly accessible. However, by leveraging the connection to entropic OT, one can construct conditional paths (e.g., Brownian bridges) between samples $(x_{0},x_{1})$ from the entropic OT plan and use these to derive analytic expressions for $u_{t}^{\circ}(x|x_{0},x_{1})$ and $\nabla\log p_{t}(x|x_{0},x_{1})$. Specifically, given a source-target pair $(x_{0},x_{1})$ sampled from the entropic OT plan $\pi^{*}_{\varepsilon}$, the conditional distribution at time $t$ is a Gaussian $p_{t}(x\mid x_{0},x_{1})=\mathcal{N}(x;(1-t)x_{0}+tx_{1},\sigma^{2}t(1-t)I)$ with constant diffusion rate $g(t)=\sigma$. From this, we obtain closed-form expressions for the conditional score and probability flow drift:
$u_{t}^{\circ}(x|x_{0},x_{1})=\frac{1-2t}{1-t}(x-((1-t)x_{0}+tx_{1}))+(x_{1}-x_{0}),\ \nabla\log p_{t}(x|x_{0},x_{1})=\frac{(1-t)x_{0}+tx_{1}-x}{\sigma^{2}t(1-t)}.$
The resulting conditional loss function combines flow and score matching terms:
$\mathcal{L}_{\mathrm{[SF]}^{2}\mathrm{M}}(\theta)=\mathop{\mathbb{E}}\limits_{\begin{subarray}{c}t\sim\mathcal{U}(0,1)\\ x\sim p_{t}(x|x_{0},x_{1})\\ (x_{0},x_{1})\sim\pi^{*}\end{subarray}}\left[\left\lVert v_{\theta}(t,x)-u_{t}^{\circ}(x|x_{0},x_{1})\right\rVert^{2}+\lambda(t)^{2}\left\lVert s_{\theta}(t,x)-\nabla\log p_{t}(x|x_{0},x_{1})\right\rVert^{2}\right],$
where $\lambda(\cdot)$ is some hyperparameter choice of positive weight. We direct readers to the appendix for more details on this formulation.
3.3 Problem Definition
We formalize the vision-based dexterous grasp translation problem as a stochastic transport task between grasp distributions associated with different robotic hands. Let $\mathcal{G}_{\mathrm{source}}\subset\mathbb{R}^{n}$ and $\mathcal{G}_{\mathrm{target}}\subset\mathbb{R}^{m}$ denote the source and target hand configuration spaces with $n$ and $m$ degrees of freedom, including the pose of the hand base in $SE(3)$, and let $\mathcal{O}$ represent the space of grasp observations. Given a segmented visual observation $(o_{\mathrm{obj}},o_{\mathrm{hand}})\in\mathcal{O}$ depicting the source hand with configuration $g_{\mathrm{source}}\in\mathcal{G}_{\mathrm{source}}$ grasping an object, our goal is to generate a corresponding grasp configuration $g_{\mathrm{target}}\in\mathcal{G}_{\mathrm{target}}$ for the target hand such that certain physical grasp properties are preserved.
We assume that grasps from the source hand follow a distribution [^1] Here, $g_{\mathrm{source}}$ is implicitly defined by $o_{\mathrm{hand}}$. $q_{0}(g_{\mathrm{source}}|o_{\mathrm{obj}})$, while feasible grasps for the target hand follow $q_{1}(g_{\mathrm{target}}|o_{\mathrm{obj}})$, both conditioned on the same object observation. These distributions capture the variability and physical plausibility of grasps within each hand’s morphology. The problem is to learn a mapping between $q_{0}$ and $q_{1}$ that respects both the structure of the source grasp and the constraints of the target hand.
To achieve this, we define a stochastic process $\{g_{\mathrm{target}}^{t}\}_{t\in[0,1]}$ interpolating between $q_{0}$ and $q_{1}$, governed by a Schrödinger Bridge with a suitable reference dynamics. The transport should minimize the deviation from a reference process (e.g., Brownian motion) while aligning the marginal distributions at $t=0$ and $t=1$. Crucially, a common choice [78, 79] in related works is to define the transport cost $d(\cdot,\cdot)$ in terms of Euclidean or geodesic distances, which primarily capture geometric dissimilarities. However, such measures are insufficient in our context, as they do not reflect the functional equivalence of grasps. Therefore, we also require a physically grounded notion of cost that considers the grasp’s effect on the object, motivating the design of more expressive transport mechanisms aligned with the task’s semantics.
The translation task thus reduces to learning a generative model for $q_{1}(g_{\mathrm{target}}|o_{\mathrm{obj}})$ that aligns with $q_{0}(g_{\mathrm{source}}|o_{\mathrm{obj}})$ under the SB formulation, leveraging both vision-based cues and physical constraints.
4 Methodology
Figure 2: Architecture overview. Blue modules correspond to stage 1: the source hand observation is encoded via a VAE. Orange modules correspond to stage 2: the latent is translated using a U-ViT Schrödinger Bridge model conditioned on object shape and contact anchors. The translated latent is decoded to produce the target hand grasp.
4.1 Latent Score and Flow Matching
We approach grasp translation as stochastic transport in a learned latent space where the high-dimensional grasp configurations for source and target hands are first embedded into a shared latent representation. This is achieved through a two-stage pipeline. In the first stage, we train a variational autoencoder (VAE) [38] that encodes segmented visual observations of the source hand grasping an object into a latent variable $z\in\mathcal{Z}$. The VAE consists of an encoder mapping $\mathcal{O}\rightarrow\mathcal{Z}$ and a decoder mapping $\mathcal{Z}\rightarrow\mathcal{G}$. The decoded grasp configuration $\hat{g}=\mathrm{Dec}(z)$ is passed through a differentiable kinematic layer $\mathrm{FK}:\mathcal{G}\rightarrow\mathbb{R}^{3N}$, yielding the 3D mesh vertices of the hand $\hat{v}=\mathrm{FK}(\hat{g})$. The VAE is optimized with a combined loss with a small KL term:
$\mathcal{L}_{VAE}=\mathbb{E}_{o\sim\mathcal{O}}[\left\lVert\hat{g}-g\right\rVert^{2}+\alpha\left\lVert\hat{v}-\mathrm{FK}(g)\right\rVert^{2}]+\beta\mathrm{KL}\left(q_{\mathrm{Enc}}(z|o)\|\mathcal{N}(0,I)\right),$ (8)
where $g$ denotes the ground truth grasp configuration and $\alpha,\beta$ are weighting coefficients. This training encourages the latent space $\mathcal{Z}$ to encode semantically meaningful and physically grounded grasp features that generalize across hands.
In the second stage, we model a Schrödinger Bridge between the latent source and target distributions, $q_{0}(z_{\mathrm{source}}|o_{\mathrm{obj}})$ and $q_{1}(z_{\mathrm{target}}|o_{\mathrm{obj}})$, with a simulation-free framework. The stochastic path $z^{t}\in\mathcal{Z}$, $t\in[0,1]$ is governed by the SDE in Eq. 3 with marginals $p_{t}(z)$ and corresponding deterministic flow ODE as in Eq. 5.
Following Tong et al. [78], we train neural networks $v_{\theta}(t,z)$ and $s_{\theta}(t,z)$ to approximate the conditional flow $u_{t}^{\circ}(z|z_{0},z_{1})$ and score $\nabla\log p_{t}(z|z_{0},z_{1})$, respectively. The training is performed using conditional Brownian bridge samples $z^{t}\sim p_{t}(z|z_{0},z_{1})$ where $(z_{0},z_{1})\sim\pi^{*}_{\varepsilon}$, which is the entropic OT plan defined in Eq. 2 between $q_{0}$ and $q_{1}$. The training objective is then defined as Eq. 7.
During inference, given a visual observation of a grasp with the source hand $(o_{\mathrm{obj}},o_{\mathrm{hand}})$, we sample a latent trajectory starting from a source encoding $z_{0}=\mathrm{Enc}(o_{\mathrm{hand}})\sim q_{0}(z_{\mathrm{source}}|o_{\mathrm{obj}})$ and evolve it forward using the learned SDE dynamics to obtain a translated grasp latent code $z_{1}\sim q_{1}(z_{\mathrm{target}}|o_{\mathrm{obj}})$. This code is decoded to produce the target hand configuration $g_{\mathrm{target}}=\mathrm{Dec(z_{1})}$, which is then used to actuate the target hand. The training and inference processes are detailed in Alg. 1, 2, and 3.
Algorithm 1 VAE Training
1:a dataset of $(o,g)$, initial network $\mathrm{Enc}$, $\mathrm{Dec}$, FK layer
2:for all $(o,g)$ in dataset do
3:     $z\sim\mathrm{Enc}(o)$
4:     $\hat{g}\leftarrow\mathrm{Dec}(z)$
5:     $\hat{v}\leftarrow\mathrm{FK}(\hat{g})$
6:     Update $\mathcal{L}_{\text{VAE}}$ with Eq. 8
7:     Update $\mathrm{Enc}$, $\mathrm{Dec}$
8:end for
9:return $\mathrm{Enc,Dec}$
Algorithm 2 Latent SB
1:Trained $\mathrm{Enc}$, source and target dataset, OTPlan, initial $v_{\theta}$,$s_{\theta}$
2:while Training do
3:    Sample $(o_{s},g_{s})$, $(o_{t},g_{t})$ from dataset
4:    $z_{s}\leftarrow\mathrm{Enc}(o_{s})$
5:    $z_{t}\leftarrow\mathrm{Enc}(o_{t})$
6:    $\pi^{*}_{\varepsilon}\leftarrow\mathrm{OTPlan}(z_{s},z_{t})$
7:    $(z_{0},z_{1})\sim\pi^{*}_{\varepsilon}$
8:    $t\sim\mathcal{U}(0,1)$
9:    $z^{t}\sim p_{t}(z\mid z_{0},z_{1})$
10:    Update $\mathcal{L}$ with Eq. 7
11:    Update $v_{\theta}$, $s_{\theta}$
12:end while
13:return $v_{\theta}$,$s_{\theta}$
Algorithm 3 Inference
1:Trained $\mathrm{Enc}$, $\mathrm{Dec}$, $v_{\theta}$, $s_{\theta}$, test observation $o_{s}$
2:$z_{0}\leftarrow\mathrm{Enc}(o_{s})$
3:$z_{1}\sim p_{1}(z|o_{s})$ using $v_{\theta}$, $s_{\theta}$
4:$\hat{g}_{t}\leftarrow\mathrm{Dec}(z_{1})$
5:return $\hat{g}_{t}$
Algorithm 1: Training and inference procedures. Left: VAE training. Center: Simulation-free Schrödinger Bridge training in latent space. Right: Inference by latent evolution.
A central challenge in learning this SB is the absence of paired samples between source and target hand configurations. Without a principled coupling, arbitrary samples from $q_{0}$ and $q_{1}$ may lead to latent translations that are uninformative or misaligned. This motivates the next section, where we introduce physically grounded ground costs for entropic OT to induce meaningful correspondences between source and target grasps.
4.2 Grasp-specific OT Cost Design
To enable meaningful grasp translation between different hand morphologies, it is crucial to guide the Schrödinger Bridge process with a ground cost that reflects task-relevant physical properties rather than naive geometric alignment. Conventional entropic OT approaches [78, 79] typically define the ground cost $d(\cdot,\cdot)$ using Euclidean or geodesic distances in either configuration or latent space. However, such metrics fail to capture the functional equivalence of grasps, particularly when source and target hands differ significantly in kinematics and contact modalities. We propose a set of physics-informed cost functions that better encode grasp quality and intent.
Base Pose Cost. We define a Euclidean cost $d_{\mathrm{pose}}$ in $\mathbb{R}^{9}$ between the global poses of the source and target hand. The pose is represented as a concatenation of the 3D base position and a 6D continuous representation of the rotation [94] to ensure smoothness and avoid ambiguities. This cost encourages coarse alignment of the grasps in workspace coordinates:
$d_{\mathrm{pose}}=\left\lVert h_{\mathrm{source}}-h_{\mathrm{target}}\right\rVert^{2}+\left\lVert R_{\mathrm{source}}-R_{\mathrm{target}}\right\rVert^{2},$ (9)
where $h_{\mathrm{source}}$, $h_{\mathrm{target}}$, $R_{\mathrm{source}}$, $R_{\mathrm{target}}$ are translation and rotation of the base pose extracted from $g_{\mathrm{source}}$, $g_{\mathrm{target}}$, respectively.
Contact Map Similarity. Each grasp is associated with a contact map indicating the location of contacts on the object surface. We represent these contact maps as 3D point clouds and compute the bidirectional Chamfer distance between the source and target maps $C_{\mathrm{source}}$, $C_{\mathrm{target}}$:
$d_{\mathrm{contact}}=\mathrm{Chamfer}(C_{\mathrm{source}},C_{\mathrm{target}})=\sum_{x\in C_{\mathrm{source}}}\min_{y\in C_{\mathrm{target}}}\left\lVert x-y\right\rVert^{2}+\sum_{y\in C_{\mathrm{target}}}\min_{x\in C_{\mathrm{source}}}\left\lVert y-x\right\rVert^{2}.$ (10)
This encourages preservation of local interaction geometry across morphologically distinct hands.
Grasp Wrench Space Overlap. The grasp wrench space (GWS) [26] characterizes the set of wrenches (forces and torques) that can be exerted on an object by a grasp. For each grasp, we compute a set of wrench vectors from contact locations and normals, then take the convex hull to approximate the GWS in $\mathbb{R}^{6}$ to form the grasp wrench hull (GWH). We define the cost based on the intersection-over-union (IoU) of GWH volumes between source and target hulls:
$d_{\mathrm{wrench}}=1-\frac{\mathrm{Vol}\left(\mathrm{Hull}({GWS}_{\mathrm{source}})\cap\mathrm{Hull}({GWS}_{\mathrm{target}})\right)}{\mathrm{Vol}\left(\mathrm{Hull}({GWS}_{\mathrm{source}})\cup\mathrm{Hull}({GWS}_{\mathrm{target}})\right)}.$ (11)
This term promotes similarity in the mechanical capabilities of the two grasps.
Jacobian-based Manipulability. To capture the grasp’s potential for object actuation, we execute the grasp in a differentiable simulator [52] and compute the Jacobian $J\in\mathbb{R}^{6\times(n-6)}$ of the object’s translational and rotational motion with respect to the hand’s joint angles, excluding the translation and rotation of the hand base. Taking a max over the joint dimension yields a 6D vector summarizing the maximal motion in each direction $m=\max_{j}|J_{j}|\in\mathbb{R}^{6}$. The cost is then defined as the squared L2 distance between these maximal effect vectors:
$d_{\mathrm{jac}}=\left\lVert m_{\mathrm{source}}-m_{\mathrm{target}}\right\rVert^{2}.$ (12)
This metric encourages the translated grasp to retain similar object-level controllability.
Minibatch Approximation. For each of the four cost functions above, we compute a separate entropic OT plan $\pi_{\varepsilon}^{*}$ using Sinkhorn’s algorithm [75]. Following Tong et al. [78], we adopt a simulation-free approach to train the SB dynamics using these plans as static couplings. Since exact OT scales quadratically with sample size, we employ minibatch entropic OT [23, 25] to efficiently estimate the plans during training. Specifically, for each minibatch of source-target samples, we compute a local OT plan using the corresponding cost matrix and regularization parameter $\varepsilon$. These minibatch couplings serve as conditioning for generating Brownian bridge paths and estimating regression targets for score and flow matching. This strategy preserves the SB formulation’s structure while enabling scalable learning in high-dimensional latent spaces.
4.3 Implementation Details
Our VAE is implemented using a PVCNN-based backbone [47]. For object shape representation, we encode the object point cloud using a LION VAE [83] pretrained on ShapeNet [12], which yields one global shape token and 512 local tokens. The hand observation latent is encoded into a single token. We further incorporate five contact anchor tokens, following the strategy introduced in [48], to guide translation via physical grounding. The latent Schrödinger Bridge is modeled using a U-ViT model [4], which serves as the backbone for $s_{\theta}$ and $v_{\theta}$.
For the Grasp Wrench Space Overlap cost, we reduce the 6D convex hull to its first three spatial dimensions, approximating the set of achievable center-of-mass forces. Monte Carlo estimation is used to evaluate the IoU of these 3D hulls efficiently on GPUs. For the Jacobian-based maximal effect descriptor, we leverage the Warp differentiable simulator [52] to simulate the grasp and compute the Jacobian of object pose with respect to joint angles. At inference time, latent samples are evolved using Euler–Maruyama integration of the learned SDE with discretized steps. No test-time finetuning is performed. Additional architectural and hyperparameter details are provided in the supplementary materials.
5 Experiments
5.1 Experimental Setup
[…]
Method Success Rate (%$\uparrow$ Diversity (rad)$\uparrow$ 6D GWH IoU (%$\uparrow$
[…]
RFP [37] 66.80 33.89 70.31 57.00 0.225 0.186 0.199 0.203 7.51 6.62 8.99 7.77
Diffusion 72.57 38.74 71.19 60.83 0.301 0.206 0.307 0.271 9.54 9.12 10.07 9.58
[…]
Dataset. We evaluate our method on the MultiGripperGrasp dataset [11], a large-scale benchmark containing 30.4 million grasps across 11 robotic manipulators and 345 objects. For our experiments, we select 138 objects for training and 34 unseen objects for testing. Our primary evaluation scenarios focus on translating grasps between an articulated human hand (5 fingers, 20 DoFs), the Allegro hand (4 fingers, 16 DoFs), and the Shadow hand (5 fingers, 22 DoFs).
Metrics. We assess performance using several quantitative metrics. First, we report the grasp success rate using IsaacGym [53], following the GenDexGrasp [42] evaluation protocol. A grasp is considered successful if it maintains a stable hold on the object under six perturbation trials, where external forces are applied along ±$\pm x$, ±$\pm y$, and ±$\pm z$ axes. We apply each force as a uniform acceleration of $0.5m/s^{2}$ for 60 simulation steps and measure whether the object translates more than 2 cm. A grasp passes if it withstands all six trials. Simulation parameters include a friction coefficient of 10 and an object density of 10,000. We use IsaacGym’s built-in positional controller to achieve the desired joint configurations. In addition to the success rate, we report the diversity of the generated grasps, computed as the average standard deviation across all revolute joints, as well as the functional similarity between source and translated grasps. The latter is measured using the full 6D IoU of GWH as defined in Eq. 11.
Baselines. We evaluate all four variants of our model, each trained using a distinct physics-based OT cost introduced in Sec 4.2. As baselines, we compare against RobotFingerPrint [37], which is a recent grasp transfer method using a unified gripper coordinate space, and a diffusion-based model trained with Contact Map Similarity OT, conditioned on the source contact map.
5.2 Quality of Translated Grasps
[…]
Among all methods, our approach using the Jacobian-based OT cost achieves the highest mean success rate and outperforms other baselines by a substantial margin. Notably, the Jacobian variant also shows consistent improvements across all tasks, demonstrating its ability to preserve functionally effective contact configurations despite differences in morphology. The GWH-based model also performs competitively, which suggests that both wrench- and control-aligned transport costs are particularly effective for grasp transfer. Qualitative results in Fig. 3 (left) also show that our method can generate stable grasps across variants, even when the source grasp is sub-optimal. In contrast, object-agnostic methods like RFP may fail to form valid grasps due to a lack of fine-grained contact reasoning.
Figure 3: (left) Qualitative results. Our method generates stable and consistent grasps across different OT variants, even when the source grasp is sub-optimal where RFP [37] fails. (right) Failure Modes where our method struggles with thin-shell objects due to challenging geometry.
On the diversity metric, the Contact OT model achieves the highest average joint-space variation. This highlights that our methods can generate diverse grasp configurations while maintaining physical feasibility. Interestingly, other OT costs show lower diversity despite achieving high success rates, indicating that geometric alignment alone may lead to more conservative grasps.
5.3 Physical Alignment between Translated Grasps
[…]
5.4 Ablation Study
[…]
$\sigma$ Success %↑ Div. (rad)↑ IoU↑
0.01 77.16 0.284 16.72
0.1 74.78 0.294 15.89
1.0 64.39 0.378 13.83
[…]
Ours Diff. Ours Diff. Ours Diff.
10 71.45 60.00 0.278 0.344 13.80 6.19
100 75.12 72.57 0.283 0.301 14.00 9.54
200 74.78 73.17 0.294 0.297 15.89 9.66
[…]
Effect of Diffusion Rate. We vary the diffusion rate $\sigma\in\{0.01,0.1,1.0\}$ and observe that smaller diffusion magnitudes lead to better grasp success and GWH IoU. Larger diffusion ($\sigma=1.0$) produces noisier paths and degrades both physical precision and grasp quality, despite yielding higher diversity. This supports the intuition that overly noisy reference processes can disrupt fine-grained transport and hinder alignment with physical intent.
Effect of Discretization Steps. We evaluate our method under varying numbers of time steps during test-time integration: {10, 100, 200}. As shown in Table 3, using 100 steps offers a strong tradeoff between accuracy and efficiency, while more steps lead to better physical alignment. Interestingly, our method remains robust even with as few as 10 steps, outperforming the diffusion baseline across all metrics and steps. In contrast, the diffusion model’s performance degrades substantially under low discretization, highlighting the benefit of our SB-based formulation, which yields more stable and sample-efficient generation dynamics.
6 Discussion
In this work, we present a novel formulation of vision-based dexterous grasp translation as a Schrödinger Bridge problem between grasp distributions. By leveraging a latent score and flow matching framework and designing physics-informed ground costs, our method enables simulation-free, distribution-level grasp transfer across robotic hands with distinct morphologies. Empirical evaluations across several hand-object settings demonstrate that our approach not only produces stable and diverse grasps but also preserves functional grasp properties more effectively than object-agnostic or geometry-based baselines. These results highlight the promise of distributional transport techniques for advancing semantic grasp understanding and generalization in robotic manipulation.
Limitations. Despite its strengths, our approach has some limitations. Most notably, as shown in Fig. 3 (right), the performance drops in scenarios involving thin-shell objects with ambiguous or minimal contact regions, where the valid grasp configuration space is largely confined. Additionally, we observe that the results for the Shadow hand are consistently lower than for the Allegro hand across multiple metrics, which aligns with prior findings in the MultiGripperGrasp dataset [11] due to its more complex kinematics and larger workspace. These discrepancies suggest that improving dataset quality and coverage for specific hands may further enhance generalization and translation fidelity.
Societal Impact. Finally, while our framework is designed for general-purpose manipulation, it introduces broader implications for deployment in real-world systems. Effective grasp translation could accelerate learning from heterogeneous demonstrations or expand manipulation capabilities in assistive robotics. However, it also raises considerations around safe deployment, failure detection, and responsible generalization across users and hardware. We encourage future work to examine how distributional grasp priors can be integrated into transparent and human-aligned manipulation pipelines.
Appendix A Implementation Details of the Overall Objectives
In our method, we adopt the [SF]2[78] training framework to jointly learn the score function $s_{\theta}(t,x)$ and the drift field $v_{\theta}(t,x)$ of the corresponding SDE and its probability flow ODE. We follow the standard formulation of the [SF]2M loss:
$\mathcal{L}_{\mathrm{[SF]}^{2}\mathrm{M}}(\theta)=\mathop{\mathbb{E}}\limits_{\begin{subarray}{c}t\sim\mathcal{U}(0,1)\\ x\sim p_{t}(x|x_{0},x_{1})\\ (x_{0},x_{1})\sim\pi^{*}\end{subarray}}\left[\left\lVert v_{\theta}(t,x)-u_{t}^{\circ}(x|x_{0},x_{1})\right\rVert^{2}+\lambda(t)^{2}\left\lVert s_{\theta}(t,x)-\nabla\log p_{t}(x|x_{0},x_{1})\right\rVert^{2}\right].$
A.1 Weighting Schedule $\lambda(t)$
Following Tong et al. [78], we apply a time-dependent weighting schedule $\lambda(t)$ to stabilize the score regression loss across time. This is essential because the conditional score $\nabla_{x}\log p_{t}(x\mid x_{0},x_{1})$ grows unbounded as $t\rightarrow 0$ or $t\rightarrow 1$, leading to imbalance in the loss. To address this, we adopt the formulation that standardizes the regression target to have unit variance. This is done by predicting the noise added in sampling $x_{t}$ from the linear interpolation $\mu_{t}=(1-t)x_{0}+tx_{1}$ under a Brownian bridge with variance $\sigma_{t}^{2}=\sigma^{2}t(1-t)$. This leads to:
$\lambda(t)=\sigma_{t}=\sigma\sqrt{t(1-t)}.$
This setting ensures that the target score $\nabla_{x}\log p_{t}(x\mid x_{0},x_{1})$ corresponds to $\epsilon/\sigma_{t}$, with $\epsilon\sim\mathcal{N}(0,I)$, making the scaled target distributed as standard Gaussian.
In practice, we use an alternative numerically stable formulation derived by rescaling the target by $\frac{1}{2}\sigma_{t}^{2}\nabla_{x}p_{t}$, which yields the weighting schedule:
$\lambda(t)=\frac{2}{\sigma^{2}}\sigma_{t}=\frac{2\sqrt{t(1-t)}}{\sigma}.$ (13)
Under this form, the squared score loss term simplifies to:
$\lambda(t)^{2}\left\lVert s_{\theta}(t,x)-\nabla\log p_{t}(x|x_{0},x_{1})\right\rVert^{2}=\left\lVert\lambda(t)s_{\theta}(t,x)+\epsilon_{t}\right\rVert.$
This avoids any division by small values resulting from boundary conditions and improves numerical stability. All models in our experiments were trained using this second formulation and the simplified regression objective:
$\mathcal{L}(\theta)=\mathop{\mathbb{E}}\limits_{\begin{subarray}{c}t\sim\mathcal{U}(0,1)\\ x\sim p_{t}(x|x_{0},x_{1})\\ (x_{0},x_{1})\sim\pi^{*}\end{subarray}}\left[\left\lVert v_{\theta}(t,x)-u_{t}^{\circ}(x|x_{0},x_{1})\right\rVert^{2}+\left\lVert\lambda(t)s_{\theta}(t,x)+\epsilon_{t}\right\rVert^{2}\right].$ (14)
Appendix B Architecture and Training Details
B.1 VAE Training
Our VAE consists of two main components: a point cloud encoder for the segmented hand point cloud, and an MLP-based latent autoencoder that encodes and decodes grasp configurations.
Hand Encoder. We use a PVCNN-based architecture [47] to process the input hand point cloud. The input to the encoder is a point cloud of size $N\times 3$ The PVCNN follows a hierarchical structure with point and voxel convolution blocks as described below:
Table 4: PVCNN Point Feature Blocks
Output Channels Num Layers Kernel Size Voxel Resolution Multiplier
64 1 32
64 2 16
128 1 16
1024 1 N/A $\times$
The final 1024-dimensional global features are passed through an MLP consisting of [256, 128] hidden units, followed by max pooling to produce a single global feature vector for each input.
Grasp Configuration VAE. The latent grasp representation is modeled using a fully connected variational autoencoder. The encoder receives the hand representation and maps it into a latent space of size 768. The decoder reconstructs the full hand joint configuration. Layer sizes are as follows:
Table 5: Latent VAE Layer Sizes
Network Input Hidden Layers Latent Dim Output Dim
Encoder 1472 [2048, 1024] 768
Decoder 768 [2048, 1024, 256] 9 + # DoFs
Training Configuration. We train the VAE for 18 epochs using a batch size of 256, Adam optimizer with a learning rate of $3\times 10^{-4}$, and 16 dataloader workers. The loss function includes a weighted combination of reconstruction losses with $\alpha=0.6$ and a KL divergence term with a very small weight ($\beta=1\times 10^{-5}$). The training time is approximately one GPU day on a single NVIDIA A6000 GPU.
B.2 Score and Flow Model Architecture
Both the score network $s_{\theta}$ and the flow network $v_{\theta}$ in our pipeline share the same U-ViT-based backbone architecture [4]. The input to each model consists of a concatenation of latent hand representations, object shape encodings, and contact anchor tokens, following the conditioning design described in Section 4.3 of the main paper.
U-ViT Backbone. We use a vision transformer architecture adapted for point-based data. The network is composed of 12 transformer blocks, each with multi-head self-attention and MLP layers. Time conditioning is injected via learned embeddings scaled by a temperature factor. The configuration is detailed in Table 6.
Table 6: U-ViT Backbone Configuration
Parameter Value
Number of Transformer Layers (Depth) 12
Embedding Dimension 512
Number of Attention Heads 8
MLP Expansion Ratio 4
Time Embedding Scale 1000.0
Dropout Rate 0.0
Attention Dropout Rate 0.1
Input Dimensions. The model is conditioned on several input tokens:
  • Latent Hand Token: 768-dimensional vector produced by the VAE encoder.
  • Global Object Token: 128-dimensional latent from the pretrained LION object encoder.
  • Local Object Tokens: 2048 points, each with 4 dimensions (XYZ + latent code).
  • Contact Anchor Tokens: 5 tokens computed by applying farthest point sampling to object points identified via thresholding (within 0.005 m) to the nearest point to the hand point cloud.
Training Configuration. We train both the score and flow networks using the loss in Eq. 14 with the weighting schedule described in Appendix A.1. The training setup is summarized in Table 7. The training time is approximately 650 GPU hours on NVIDIA L40 GPUs.
Table 7: Training Hyperparameters for Score and Flow Models
Hyperparameter Value
Total Steps 20000
Batch Size (on each GPU) 512
Learning Rate $2\times 10^{-4}$
Linear Warmup Steps 256
Gradient Clipping 1.0
EMA Decay 0.999
EMA Start Step 10000
$\sigma$ (used in Table. 1) 0.1
Appendix C OT Cost Formulation and Computation
We implement four distinct ground costs for the optimal transport (OT) coupling in our framework. Two of these (Euclidean base pose and contact map Chamfer distance) are computed in closed form using standard metrics in $\mathbb{R}^{n}$. The remaining two involve more complex geometry- and physics-informed reasoning.
Grasp Wrench Space Overlap. The grasp wrench space (GWS) [26] of a grasp represents the set of all wrenches (forces and torques in $\mathbb{R}^{6}$) that the grasp can apply to an object through contact. For a grasp with $k$ contact points $\{c_{i}\}_{i=1}^{k}$ and associated contact normals $\{n_{i}\}$, we construct a set of unit wrench vectors $\{w_{i}\}$ as follows:
$w_{i}=\begin{bmatrix}f_{i}\\ c_{i}\times f_{i}\end{bmatrix}\in\mathbb{R}^{6},\quad f_{i}=\alpha_{i}n_{i},\ \alpha_{i}>0.$
We form the convex hull of these wrenches to approximate the GWS, denoted as the grasp wrench hull (GWH). The distance between two grasps is then defined as the 1 minus the intersection-over-union (IoU) of their GWH volumes:
$d_{\mathrm{wrench}}=1-\frac{\mathrm{Vol}(\mathrm{Hull}(GWS_{\mathrm{source}})\cap\mathrm{Hull}(GWS_{\mathrm{target}}))}{\mathrm{Vol}(\mathrm{Hull}(GWS_{\mathrm{source}})\cup\mathrm{Hull}(GWS_{\mathrm{target}}))}.$
Since computing the exact intersection and union volumes of 6D convex hulls is intractable in closed form, we approximate the GWH IoU using GPU-accelerated Monte Carlo sampling. We sample a large number of points (typically $O(10^{6})$) in the 6D bounding box enclosing both hulls, and compute membership in each hull using their supporting hyperplane inequalities. Intersection and union volumes are then estimated as:
$\text{Vol}_{\text{intersect}}\approx\frac{\#\text{pts in both}}{\#\text{samples}}\text{Vol}_{\text{box}}.$
A 3D implementation of this approximation is used to compute batch-wise GWH IoU efficiently during training, and the full 6D implementation is reported for the experiments.
Jacobian-based Manipulability. To capture the potential of a grasp to actuate the object, we compute the Jacobian matrix $J\in\mathbb{R}^{6\times(n-6)}$ relating hand joint velocities to object motion. The Jacobian is computed using the Warp differentiable simulator [52] following the setup from Turpin et al. [81] with fixed object and grasp configurations. We exclude the 6-DoF base pose of the hand, focusing only on internal joints.
Each column $J_{j}$ of the Jacobian represents the effect of joint $j$ on the object’s linear and angular velocity. We summarize the object-level actuation capability of a grasp by computing a 6D vector $m$ of maximal manipulability across joints:
$m=\max_{j}|J_{j}|\in\mathbb{R}^{6}.$
The OT cost is then defined as the squared $\ell_{2}$ distance between source and target maximal actuation vectors:
$d_{\mathrm{jac}}=\|m_{\mathrm{source}}-m_{\mathrm{target}}\|^{2}.$
This cost encourages the translated grasp to maintain similar controllability over object motion.
Appendix D Additional Results
Figure 4: More qualitative examples. Each column shows a source grasp and its translated output for different methods. Our method consistently recovers physically plausible grasps across tasks, while RFP frequently fails, especially with noisy source grasps.
D.1 Qualitative Examples Across Tasks
[…]
D.2 Inference Time Comparison
We compare the inference time of our method against both a diffusion-based baseline and the optimization-based RobotFingerPrint [37] method. As shown in Table 8, our model supports fully amortized generation and remains efficient at inference time, taking approximately 0.8 seconds per grasp using 100 Euler–Maruyama steps. This is comparable to the diffusion baseline (0.5s) and significantly faster than RFP, which requires costly per-grasp optimization.
Table 8: Average inference time per grasp (amortized).
Method Time (sec/grasp)
Diffusion (100 steps) 0.5
Ours (100 steps) 0.8
RFP [37] 5.2
Appendix E Compute Resources
Our experiments were conducted on a combination of local servers and a high-performance computing cluster. The local server consists of a 24-core CPU and 2 NVIDIA A6000 GPUs, which were used for model development, ablation studies, and VAE training. For large-scale training, we utilized a compute cluster, where each cluster node is equipped with two 26-core CPUs and 8 NVIDIA L40 GPUs.